OpenStudy (anonymous):

((x+3)/(30x-42))*((40x-56)/(8))

OpenStudy (anonymous):

\[\frac{ x+3 }{ 30x-42 }*\frac{ 40x-56 }{ 8 }\]

Nnesha (nnesha):

mango > watermelon

OpenStudy (adi3):

two of the great scientist's pictures!!!!!!!!

OpenStudy (adi3):

@Michele_Laino @IrishBoy123

OpenStudy (adi3):

what a pic

OpenStudy (adi3):

And the thing is that they both are good at mahts!!!!!!!

OpenStudy (adi3):

sorry

OpenStudy (michele_laino):

thanks! :) @Adi3

OpenStudy (adi3):

np

OpenStudy (adi3):

try cross multiply

OpenStudy (michele_laino):

@helpme375 please note that your expression can berewritten as follows: \[\frac{{x + 3}}{{6\left( {5x - 7} \right)}} \cdot \frac{{8\left( {5x - 7} \right)}}{8}\]

OpenStudy (adi3):

try cross multiply

Nnesha (nnesha):

irish was typing so i thought he is going to help you.

OpenStudy (michele_laino):

I think that answer is: \[\frac{{x + 3}}{6}\]

OpenStudy (anonymous):

how did you get it

OpenStudy (michele_laino):

when you simplify that expression above, then factors 5x-7 cancel, also factors 8 cancel, so we reamin with (x+3) as numerator and 6 as denominator

OpenStudy (michele_laino):

remain*

OpenStudy (anonymous):

but would you be left with (x+3)/8 after you take away 5x-7

OpenStudy (michele_laino):

here is the right simplification: 

OpenStudy (anonymous):

okay thanks i got it

OpenStudy (michele_laino):

:)